Text Size

April 11, 2010
ISS On-Orbit Status 04/11/10

All ISS systems continue to function nominally, except those noted previously or below. FD7 (Flight Day 7) of STS-131/19A. Crew sleep shifting in effect – see below. Sunday – full work day. Ahead: Week 4 of Increment 23.

Mission 19A’s EVA-2 was completed by EV1 Rick Mastracchio & EV2 Clayton Anderson in 7h 26m, accomplishing almost all major objectives. Beginning this morning at 1:30am EDT, the spacewalk ended at 8:58am. Due to a difficulty with bolt tightening during installation of the new ATA (Ammonia Tank Assembly) on the S1 truss, the spacewalkers fell behind on the timeline and could not complete all objectives. [EV1 & EV2 began their “campout” yesterday after noon (~12:31pm) in the U.S. Airlock (A/L) with hatch closure and depressurization of the Crewlock (CL) from 14.7 to 10.2 psi, followed by mask prebreathe. Following the usual hygiene break/with mask prebreathe for Mastracchio & Anderson at ~10:51pm, the A/L hatch was closed again by TJ Creamer & Tracy Caldwell-Dyson for EVA preps in 10.2 psi, followed by EMU purge (~12:41am) and prebreathe in the EMUs (12:56am-1:46am). Afterwards, with CL depressurization and EV1/EV2 switching to suit power, EVA-2 began at 1:30am. The excursion lasted 7h 26m.]

During EVA-2, Mastracchio & Anderson –
· Transferred the spent ATA to the CETA (Crew & Equipment Translation Aid);
· Attached the new ATA structurally (4 bolts) on the S1 truss,
· Mated electrical connectors to the new ATA to apply heater power,
· Installed P1 Radiator Grapple Fixture Stowage Beams,
· Relocated one (of two) APRF (Articulated Portable Foot Restraint) to CETA cart WIF2 (Worksite Interface 2) for the ULF4 mission,
· Stowed the old ATA on the POA (Payload ORU Accommodation).

Three objectives were deferred for EVA-3 (being replanned at this time):
  • Mating ATA fluid QDs (quick disconnects) for NH3 (ammonia) & N2 (nitrogen) on the new ATA,
  • Removing/retrieving A/L MMOD (Airlock Micrometeoroid/Orbital Debris) shields from ESP-2 (External Stowage Platform 2), and
  • Relocating the second APFR to CETA WIF1.

[Background: Engaging & torquing one of the four bolts took considerably longer than planned but was eventually accomplished. By that time, the EVA had taken 5h 15m which did not allow enough time to complete all planned tasks. Although Rick & Clay had sufficient O2 (oxygen) & Metox CO2 (carbon dioxide) filtration consumables for an 8h EVA, hooking up the fluid connectors, with the risk of NH3 spillage on the spacesuits, raised some concern that if a contamination happened, the EVA crew would not have enough time to do a suit “bake-out” in the sunlight before ingressing, thus potentially getting toxic ammonia into the A/L.]

Before the EVA, CDR Kotov verified closure of the external protective shutters of the Lab, Kibo and Cupola windows while FE-6 Creamer powered down the amateur/ham radio equipment to prevent RF interference with the spacewalkers.

FE-5 Noguchi set up the RWS (Robotic Workstation) and DOUG (Dynamic Onboard Ubiquitous Graphics) peripherals with the IPV (International Procedures Viewer) laptop to support SSRMS (Space Station Remote Manipulator System) ops.

During the spacewalk, PLT Jim Dutton & MS3 Steph Wilson operated the Canadarm-2 SSRMS (Space Station Remote Manipulator System), supporting Rick & Clay with transferring the old ATA to CETA and later to POA, plus the transfer of the fresh ATA to the S1 truss. [Ground operators later (~11:30am EDT) “walked” the SSRMS to the Node-2 PDGF (Power & Data Grapple Fixture) and maneuvered it into position to support EVA-3.]

At wake-up (9:21 pm last night), FE-1 Skvortsov performed the regular daily early-morning check of the aerosol filters at the Russian Elektron O2 generator which Maxim Suraev had installed on 10/19 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [FE-1 again inspected the filters before bedtime this noon, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

FE-3 Kornienko stowed the CDM (Carbon Dioxide Monitor, #1020) and its pump filter in CDMK (CDM Kit) 1010. [CDM was used overnight to check CO2 accumulation in the vicinity of Mikhail’s head during sleep in his new sleep station in the JAXA JLP (JEM Logistics Pressurized Segment).]

After breakfast, FE-5 Noguchi once more supported the weekly U.S. “Bisphosphonates” biomedical countermeasures experiment, ingesting an Alendronate pill before breakfast. [The Bisphosphonates study should determine whether antiresorptive agents (that help reduce bone loss) in conjunction with the routine in-flight exercise program will protect ISS crewmembers from the regional decreases in bone mineral density documented on previous ISS missions. Two dosing regimens are being tested: (1) an oral dose of 70 mg of Alendronate taken weekly starting 3 weeks prior to flight and then throughout the flight and (2) an intravenous (IV) dose of 4 mg Zoledronic Acid, administered just once approximately 45 days before flight. The rationale for including both Alendronate and Zoledronic Acid is that two dosing options will maximize crew participation, increase the countermeasure options available to flight surgeons, increase scientific opportunities, and minimize the effects of operational and logistical constraints. The primary measurement objective is to obtain preflight and postflight QCT (Quantitative Computed Tomography) scans of the hip. The QCT scans will provide volumetric bone density information of both cortical and trabecular (spongy) bone regions of the hip.]

FE-2 Caldwell-Dyson supported the spacewalkers on their pre-EVA activities as IV (Intravehicular) crewmember, including campout, hygiene break, EMU assistance, hatch closure & leak check.

During EVA-2, CDR Poindexter & MS2 Metcalf-Lindenburger (ML) provided IV support, with Dex in charge of photo/video activities.

Post-ingress activities by Anderson, Mastracchio, Creamer & Caldwell-Dyson included the usual post-EVA tasks like photographing EMU gloves for inspection, recharging EMUs with water, downloading & downlinking D2XS EVA & glove photographs, recharging REBA batteries, etc.

Besides EVA-2, the crew’s day was dominated by MPLM Leonardo and Shuttle Middeck cargo transfers:
  • FE-5 Noguchi transferred the VCA-1 (Video Camera Assembly 1) camcorder from the COL (Columbus Orbital Laboratory) to the Node-2, mounting it on the zenith “ceiling” for monitoring cargo transfer activities in the MPLM/Node-2 area. [It was returned to the COL at the end of the crew day by Tracy and installed on its seat track at the Port Endcone, aft side.]
  • Noguchi also scavenged the first GLA (General Luminaire Assembly) from the MPLM (loc. PA3) prior to the TeSS (Temporary Sleep Station) transfer, in order to avoid possible physical interference with a later scavenge,
  • MS3 Stephanie Wilson & MS4 Naoko Yamazaki transferred the large TeSS pieces from the Lab to the MPLM for return,
  • Yamazaki reconfigured the HRF2 (Human Research Facility 2) Rack by removing and returning three drawers, then installing two new stowage drawers,
  • Naoko & Soichi transferred an unpowered GLACIER (General Laboratory Active Cryogenic ISS Experiment Refrigerator) freezer from the MPLM and installed it in the Lab (loc. O1).
  • TJ Creamer transferred the KID (KUBIK Interface Drawer) into the EDR Rack, and
  • Noguchi retrieved Area Dosimeters and installed new ones in the JAXA JPM (JEM Pressurized Module) and JLP (JEM Logistics Pressurized Segment).

FE-1 Skvortsov & CDR Kotov set up equipment to capture the sound of the TVIS treadmill in operation, in an attempt to isolate the source of an anomalous noise reported by the crew yesterday. [The TVIS is currently NoGo while ground teams analyzes the noise and develops a forward plan to get TVIS back in a good configuration for exercise. Oleg played back/downlinked the recorded noise via high-definition MPC (Multipurpose Converter).]

Kotov configured the equipment for his third onboard session with the Russian experiment DYKHANIE (MBI-18, “Respiration”) and undertook the test, controlled from the RSE-Med laptop and supported by ground specialist tagup. Oleg then closed down the hardware and stowed it. [Dykhanie-1 uses two body belts (PG-T/thoracic, PG-A/abdominal), a calibrator, resistor, mouthpiece, etc., to study fundamental physiological mechanisms of the external breathing function of crewmembers under long-duration orbital flight conditions. During the experiment, physiological measurements are taken and recorded with a pneumotachogram, a thoracic pneumogram, an abdominal pneumogram, and pressure data in the oral cavity. All experimentally derived plus salient environmental data along with personal data of the subject are recorded on PCMCIA (Personal Computer Memory Card International Association) card for return to the ground at end of the Expedition. Objectives include determining the dynamics of the relationship between thoracic (pectoral) and abdominal breathing function reserves and their realization potential during spontaneous breathing, the coordinated spontaneous respiratory movements in terms of thoracic and abdominal components of volumetric, time & rate parameters of spontaneous respiratory cycle, identification of the features of humoral-reflex regulation of breathing by dynamics of ventilation sensitivity of thoracic and abdominal components to chemoreceptor stimuli, etc. Overall, the experiment is intended to provide a better understanding of the basic mechanisms of pulmonary respiration/gas exchange gravitational relations of cosmonauts.]

Timothy Creamer performed the scheduled Catalytic Reactor ORU R&R (Removal & Replacement) in WRS-1 (Water Recovery System 1) and installed a filter kit for an initial system flush following ORU (Orbit Replaceable Unit) replacement. [After the R&R, ground teams undertook an O2 system fine leak check into crew sleep.]

Working in the SM PkhO (Service Module Transfer Compartment) on the internal surface of the +Y-axis ASP-GB2 docking assembly, where Maxim Suraev had removed MLI (Multilayer Insulation) on 12/3/09 for a photographic inspection of the surface for remains of film or adhesive, Skvortsov & Kotov installed a newly delivered protective cover in place of the MLI, then took documentary photography for downlink. [The MRM2 (Mini-Research Module 2) is docked at the PkhO zenith docking port.]

Continuing the current round of periodic preventive maintenance of RS (Russian Segment) ventilation systems, Mikhail Kornienko worked in the SM and FGB (Funktsionalnyi-Grusovoi Blok). [Using a vacuum cleaner and soft brush, Misha cleaned filters and fan grilles of the Group E fan grills in the SM (VPkhO, FS5, FS6, VP), central ventilators TsV1,2, the detachable VT7 fan screens of the three SOTR gas-liquid heat exchangers (GZhT4), plus the fixed GZhT4 grill, the TsV1 circulation fan and interior closeout panel vent screens (panels 201, 301, 401), and also replaced the PS1 & PS2 dust filter cartridges. Sometimes this type of work can cause false smoke warnings, as happened yesterday, when smoke detectors were triggered by stirred up dust particles.]

Oleg & Misha completed the periodic (currently daily) checkout & performance verification of IP-1 airflow sensors in the various RS hatchways. [Inspected IP-1s are in the passageways PrK (SM Transfer Tunnel)–RO (SM Working Compartment), PkhO (SM Transfer Compartment)–RO, PkhO–DC1, PkhO–FGB PGO, PkhO-MRM2, FGB PGO–FGB GA, and FGB GA–Node-1.]

Soichi configured the VDS (Video Distribution System) television set-up in Node-2, verifying that the video cap was installed which enables pass-through reception of video from the Discovery with the Orbiter docked in support of SSRMS (Space Station Remote Manipulator System ops.

FE-5 also performed regular service on the WHC (Waste & Hygiene Compartment), changing out its UR (Urine Receptacle) hose and IF (Insert Filter), then vacuumed the entire WHC and cleaned it with disinfectant wipes. [The old UR & IF were double-bagged and stowed for disposal.]

Kotov , Skvortsov & Kornienko joined up for the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP-Moscow. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Oleg, Alexander & Mikhail also checked up on the Russian POTOK-150MK (150 micron) air filter unit of the SM’s SOGS air revitalization subsystem, gathering weekly data on total operating time & “On” durations for reporting to TsUP.

An additional task for the Russian crew in the SM was a routine inspection of the SVO SRV-K2M Condensate Water Processor hoses from the MFR Diaphragm Separator Filter to the BRPK Condensate Separation & Pumping Unit (the latter having been replaced on 12/21/09).

Skvortsov completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

To check for a possible leak, TJ Creamer marked a PBA (Portable Breathing Apparatus, #1028) in the JPM with a with a Sharpie marker at the current pressure gauge value.

The crew completed today’s physical workout regime on the CEVIS cycle ergometer (FE-5, FE-6), ARED advanced resistive exerciser (CDR, FE-2, FE-3, FE-5, FE-6), T2/COLBERT advanced treadmill (CDR, FE-1, FE-2, FE-3) and VELO ergometer bike with bungee cord load trainer (FE1).

Transfers: 8 CWC-I’s (Collapsible Water Containers-Iodine) and 2 CWCs have been filled and transferred. Cargo transfers are ahead of schedule at 71% overall complete, with 71% complete in the Middeck, and 71% complete in the MPLM.

Wake/Sleep schedule on ISS (EDT):


No CEO (Crew Earth Observations) photo targets uplinked for today.

ISS Orbit (as of this morning, 8:40am EDT [= epoch])
Mean altitude – 345.6 km
Apogee height – 348.9 km
Perigee height – 342.3 km
Period -- 91.45 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0004914
Solar Beta Angle -- -5.9 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.75
Mean altitude loss in the last 24 hours – 144 m
Revolutions since FGB/Zarya launch (Nov. 98) – 65,297

Significant Events Ahead (all dates Eastern Time and subject to change):
--------------Six-crew operations-----------------
04/07/10 -- STS-131/Discovery/19A
· 04/13 -- EVA-3 (3:11am)
· 04/15 -- MPLM install in PLB: 6:51am-10:21am
· 04/17 – FD13 – undocking: 10:25am
· 04/19 – FD15 – deorbit burn: 7:51am
04/19/10 -- STS-131/Discovery/19A – FD15 – landing: 8:53am
04/27/10 -- Progress M-03M/35P undock
04/28/10 -- Progress M-05M/37P launch
04/30/10 -- Progress M-05M/37P docking
05/10/10 -- Progress M-04M/36P undock
05/12/10 -- Soyuz 21S relocation (FGB Nadir to SM Aft)
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1 “Rassvet”
06/02/10 -- Soyuz TMA-17/21S undock/landing (End of Increment 23)
--------------Three-crew operations-------------
06/14/10 -- Soyuz TMA-19/23S launch – Wheelock (CDR-25)/Walker/Yurchikhin
06/16/10 -- Soyuz TMA-19/23S docking
--------------Six-crew operations-----------------
06/28/10 -- Progress M-06M/38P launch
06/30/10 -- Progress M-06M/38P docking
07/07/10 -- US EVA-15 (Caldwell/Wheelock)
0*7/23/10 -- Russian EVA-25 (Yurchikhin/Kornienko)
07/26/10 -- Progress M-05M/37P undock
07/29/10 -- STS-134/Endeavour (ULF6 – ELC3, AMS-02)
08/30/10 -- Progress M-06M/38P undock
08/31/10 -- Progress M-07M/39P launch
09/02/10 -- Progress M-07M/39P docking
09/16/10 -- Soyuz TMA-18/22S undock/landing (End of Increment 24)
09/16/10 -- STS-133/Discovery (ULF5 – ELC4, PMM)
09/18/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) docking
09/22/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) undock
09/30/10 -- Soyuz TMA-20/24S launch – Kelly (CDR-26)/Kaleri/Skripochka
10/xx/10 -- Russian EVA-26
10/27/10 -- Progress M-08M/40P launch
10/29/10 -- Progress M-08M/40P docking
11/26/10 -- Soyuz TMA-19/23S undock/landing (End of Increment 25)
12/10/10 -- Soyuz TMA-21/25S launch – Kondratyev (CDR-27)/Coleman/Nespoli
12/15/10 -- Progress M-07M/39P undock
12/26/10 -- Progress M-08M/40P undock
12/27/10 -- Progress M-09M/41P launch
12/29/10 -- Progress M-09M/41P docking
03/16/11 -- Soyuz TMA-20/24S undock/landing (End of Increment 26)
03/30/11 -- Soyuz TMA-22/26S launch – A. Borisienko (CDR-28)/R, Garan/A.Samokutayev
04/01/11 -- Soyuz TMA-22/26S docking
04/27/11 -- Progress M-09M/41P undock
04/28/11 -- Progress M-10M/42P launch
04/30/11 -- Progress M-10M/42P docking
05/17/11 -- Soyuz TMA-21/25S undock/landing (End of Increment 27)
05/31/11 -- Soyuz TMA-23/27S launch – M. Fossum (CDR-29)/S. Furukawa/S. Volkov
06/02/11 -- Soyuz TMA-23/27S docking
06/21/11 -- Progress M-11M/43P launch
06/23/11 -- Progress M-11M/43P docking
08/30/11 -- Progress M-12M/44P launch
09/01/11 -- Progress M-12M/44P docking
09/30/11 -- Soyuz TMA-24/28S launch
10/28/11 -- Progress M-13M/45P launch
10/30/11 -- Progress M-13M/45P docking
11/25/11 -- Soyuz TMA-25/29S launch
11/27/11 -- Soyuz TMA-25/29S docking
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton.